# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 2.8. Related works and positioning
Hyperuniformity
@@
We have computed the logarithmic derivatives of random point fields with two-dimensional Coulomb interaction potentials for the sineβ, Airyβ $(\beta=1,2,4)$, Bessel2, and Ginibre random point fields [25, 26, 35, 10]. These computations require strong hyperuniformity of particle numbers within $S_{R}$ for each $R\in\mathbb{N}$, a property that must be established separately for each random point field. For example, in the case of the Ginibre random point field, we used the following:
@@
$\sup_{N\in\mathbb{N}}\mathrm{Var}^{\mu^{N}}[N_{R}(\mathsf{s})]=O(R)\quad\text{ as }R\to\infty.$ (2.66)
@@
Here, $N_{R}(\mathsf{s})$ is the number of particles in $S_{R}$ [39, 26, 31, 32]. Such an estimate is called hyperuniform because, if $\mu^{N}$ is replaced by $\Lambda^{N}$, then the right-hand side becomes $O(R^{2})$ [20]. Here, $\Lambda^{N}$ is the conditioned Poisson random point field with Lebesgue intensity $dx$ defined as
@@
where $R(N)$ is the positive number such that the volume of $S_{R(N)}$ coincides with $N$. We also note that $\Lambda^{N}$ converges to $\Lambda$ as $N\to\infty$.
We set $\mathsf{s}=\sum_{i}\delta_{s^{i}}$. From (2.66), we have that, for each $Q\in\mathbb{N}$ and ${x}\in\mathbb{R}^{2}$,
$\lim_{R\to\infty}\limsup_{N\to\infty}\int_{S_{Q}\times\mathsf{S}}\Big{\lvert}\Big{(}\sum_{\lvert{x}-s^{i}\rvert\geq R}\frac{{x}-s^{i}}{\lvert{x}-s^{i}\rvert^{2}}\Big{)}-{x}\Big{\rvert}^{2}d\mu^{N,[1]}=0.$ (2.67)
The corrector x{x} appears in the summand in (2.67) because the measure $\mu^{N}$ lacks translation invariance and the gradient of the Coulomb potential is not integrable at infinity. From this expression, we obtained the logarithmic derivative of the Ginibre random point field [25, 34], highlighting the crucial role of hyperuniformity—similar to (2.66)—in the construction of logarithmic derivatives. While Leblé [20] established a hyperuniform estimate for general $\beta>0$, it is significantly weaker than (2.66) and insufficient for our purposes. For the sineβ, Airyβ $(\beta=1,2,4)$, Bessel2, and Ginibre random point fields, uniform estimates for the associated $N$-particle systems were obtained in [25, 26, 35, 10], relying on the specific structure of each field and techniques from random matrix theory and determinantal random point fields. However, these proofs do not extend to general Coulomb random point fields, except in the Ginibre case, necessitating a new approach.
Quenching
In this work we introduce a robust method for estimating the logarithmic derivatives of $\mu^{N}$ (Lemma 4.3) and refine it for conditional probabilities $\mu_{R,\mathsf{y}}^{N}$ (Lemma 5.2), exploiting a form of harmonicity in the interaction potential. As a central result, we obtain the uniform bound
$\sup_{N\in\mathbb{N}}\int_{S_{Q}\times\mathsf{S}}\bigl{|}\mathsf{d}^{N}({x},\mathsf{s})\bigr{|}^{2}\,\mu^{N,[1]}(d{x}d\mathsf{s})<\infty,$ (2.68)
which provides the necessary control to rigorously define logarithmic derivatives in the Coulomb setting.
The proof of (2.68) is remarkably simple: it relies only on (2.16), (2.17), and the basic definition of the logarithmic derivative. The estimate itself may be viewed as a form of hyperuniformity, since the quantity in (2.68) diverges when $\mu^{N,[1]}$ is replaced by the one-Campbell measure $\Lambda^{N,[1]}$ of the conditioned Poisson point process $\Lambda^{N}$, even after excluding a neighborhood of x{x}.
In Lemma 5.2, we establish the quenched analogue of (2.68):
$\sup_{R\geq Q+1}\sup_{N\in\mathbb{N}}\int_{S_{Q}\times\mathsf{S}}\bigl{\lvert}\mathsf{d}_{R,\mathsf{y}^{N}}^{N}({x},\mathsf{s})\bigr{\rvert}^{2}\,\mu_{R,\mathsf{y}^{N}}^{N,[1]}(d{x}d\mathsf{s})<\infty\quad\text{ for $\mu$-a.s., $\mathsf{y}$},$ (2.69)
where $\mathsf{d}_{R,\mathsf{y}^{N}}^{N}$ denotes the logarithmic derivative of $\mu_{R,\mathsf{y}^{N}}^{N}$, and $\mathsf{y}^{N}$ is defined in (5.3).
Inequality (2.69) plays a decisive role in the proof of Theorem 2.1. Indeed, it fulfills the role previously played by strong hyperuniform estimates, but with a crucial difference: its proof is significantly simpler. Earlier results relied on bounding average fluctuations in the exterior configuration, whereas here we directly control the external condition in a quenched form. This shift in perspective—replacing averaged estimates by a quenched analysis—is one of the key conceptual innovations of the present paper.
Logarithmic derivatives
The concept of the logarithmic derivative for random point fields was introduced in [25]. The present paper further develops this idea and presents the $\bullet$-logarithmic derivative (see Definition 2.2). Combining the concept of the $\bullet$-logarithmic derivative, the coupling of logarithmic derivatives in (5.5), and a quench $L^{2}$-bound of logarithmic derivatives in Lemma 5.2 and Theorem 5.1, we derive an explicit formula for the logarithmic derivative in (2.33). From (2.33), we construct the stochastic dynamics of infinite-particle systems and verify the assumptions of Theorem 2.3. Using (2.33), we prove that ISDE (2.41) has a unique strong solution (Theorem 2.3).
Until now, the logarithmic derivatives of Coulomb random point fields have been explicitly computed only in the case of the two-dimensional Coulomb field at inverse temperature $\beta=2$, namely the Ginibre random point field. These computations relied heavily on the special algebraic and determinantal structure of that particular field and do not extend to general $d$-dimensional Coulomb random point fields. In contrast, we present here a new, remarkably simple and transparent method that yields uniform estimates for the logarithmic derivative of $\mu^{N}$ (Lemma 4.3) and for its conditional analogue (Lemma 5.2), valid simultaneously for all spatial dimensions $d\geq 2$ and all inverse temperatures $\beta>0$. The key insight is to exploit a certain harmonicity property of the interaction potential, which permits effective control over the infinite-dimensional interactions. Our approach is robust, conceptually clear, and provides a versatile tool for analyzing the stochastic dynamics of $d$-dimensional Coulomb random point fields, thereby overcoming the limitations of all previously known methods.
Interacting Brownian motions
Lang initiated the systematic study of interacting Brownian motions, which was subsequently developed by Fritz, Tanemura, and others in the case of finite-range interaction potentials [17, 18, 6, 42]. When the interaction becomes long-ranged, however, the problem of constructing solutions to ISDEs poses significant difficulties, even for potentials within Ruelle’s class—namely, those that are superstable and integrable at infinity. Later, the first author introduced the Dirichlet form approach, which made it possible to establish the existence of weak solutions to ISDEs for a broad family of interaction potentials [24, 25, 26]. This framework encompasses, in particular, the Ginibre interacting Brownian motion. Nonetheless, the stronger questions of constructing pathwise unique strong solutions and establishing their uniqueness remained open.
In [46], Tsai succeeded in constructing pathwise unique strong solutions of ISDEs for the $\sin_{\beta}$ random point fields with all $\beta\geq 1$. These processes correspond to infinite-particle systems in one dimension with logarithmic interactions, that is, with the two-dimensional Coulomb potential. In this sense, the ISDE in [46] can be regarded as involving a Riesz potential. Tsai’s proof relies on a delicate coupling mechanism that exploits special properties of the logarithmic interaction in one dimension. Such an approach, however, has no counterpart in higher dimensions with genuine Coulomb potentials.
To apply this framework, one requires both an equilibrium state (random point field) and the corresponding logarithmic derivative. Moreover, a strong hyperuniform estimate of particle numbers in expanding balls of radius $R$ is typically necessary in order to control the pronounced long-range effects of Coulomb interactions. Such sharp hyperuniform estimates, however, are highly nontrivial and are known only in special cases tied to random matrix ensembles and particular inverse temperatures.
The present work introduces a new and more robust approach to circumvent this obstacle. Rather than relying on delicate hyperuniform bounds, we develop a precise representation of the logarithmic derivative (Theorem 2.1). This representation allows us to verify directly the key assumption—the IFC condition—required in the theory of [34, 16], and thereby establish the strong well-posedness of the Coulomb interacting Brownian motions for all $d\geq 2$ and $\beta>0$.
Our new approach is highly robust, and its scope is not confined to Coulomb potentials. We expect that it can be extended to Riesz potentials in arbitrary dimensions $d\geq 1$ and for all inverse temperatures $\beta>0$, both in the bulk and at the soft edge. These extensions, however, lie beyond the scope of the present work and are left for future investigation.
As already emphasized, the logarithmic derivative plays a central role in the stochastic dynamics of infinite-particle systems arising from random point fields. There exist, however, important examples of random point fields that do not admit a logarithmic derivative in the present sense. Indeed, in [37], the second author proved that for random point fields exhibiting rigidities stronger than number rigidity—such as the zero set of Gaussian analytic functions [8, 9]—a logarithmic derivative does not exist. Developing a new framework for logarithmic derivatives suitable for such highly rigid random point fields, and establishing ISDEs for them within the scheme of [34, 16], thus presents a challenging and important direction for future research.
Katori and Tanemura proposed an alternative approach to constructing stochastic dynamics with logarithmic interaction in one dimension. Their method, based on space-time correlation functions, is restricted to the special case of inverse temperature $\beta=2$ in one-dimensional systems. For further details, see [13] and the references therein. Subsequently, in [33], it was shown that the dynamics obtained through this correlation function method coincide with those described by the ISDE framework developed in [25, 34, 16].
Moreover, [33] investigated in detail the relationship between the logarithmic derivative and the space-time correlation approach. It was rigorously established that the stochastic dynamics constructed by these two seemingly distinct methods are, in fact, identical. This coincidence highlights the fundamental role of the logarithmic derivative in unifying different perspectives on infinite-particle dynamics with logarithmic interactions.
In [2], Assiotis and Mirsajjadi constructed a non-equilibrium dynamics whose equilibrium distribution is the inverse Bessel random point field, and they identified the ISDE governing this dynamics. Their method is algebraic in nature, relying on intertwining theory and the use of derivatives of the so-called “characteristic polynomial.” This construction bears a striking resemblance to the framework of logarithmic derivatives developed in [25]. We believe that clarifying the precise relationship between these two approaches is not merely of technical interest, but represents a fundamental step toward a unified understanding of stochastic dynamics associated with random point fields. In particular, such a connection could open the way to extending algebraic methods to broader classes of interactions, and conversely, to interpreting analytic techniques such as logarithmic derivatives in terms of deeper algebraic structures.
In [40], Suzuki established the ergodicity of the unlabeled dynamics of the Ginibre interacting Brownian motion, thereby providing a concrete confirmation of the general ergodicity theory (see also [36, 37]). He further derived curvature bounds for Dyson Brownian motion in [41], offering additional structural insight into these dynamics.
In the Ginibre case, the vanishing of self-diffusion was established in [28], with number rigidity of the stationary field serving as the decisive mechanism. This phenomenon connects naturally with our framework, and its broader implications for translation-invariant Coulomb systems are discussed in the Introduction. In [23], it was proved that the self-diffusion coefficient is strictly positive for any potential in Ruelle’s class with convex hardcore. Taken together, these results highlight the structural distinction between Coulomb interactions and Ruelle-type interactions.
Structure of the paper
After presenting our main theorems in this section, the remainder of the paper develops the analytic and probabilistic framework needed to establish them.
Section 3 presents concrete examples. Section 4 constructs the logarithmic derivative of Coulomb random point fields, a central analytic ingredient for the ISDE theory. Section 5 strengthens this by proving a quenched estimate and deriving an explicit representation of the logarithmic derivative.
In Section 6, we establish non-collision properties of ISDE solutions by direct SDE techniques. These arguments do not rely on Dirichlet form theory, yet the resulting non-collision property is later used to prove uniqueness of Dirichlet forms.
Sections 79 form the technical core of the paper: we introduce Dirichlet form schemes for $m$-labeled dynamics, construct infinite-volume processes under both lower and upper Dirichlet forms.
In Section 10, we verify the IFC condition ensuring consistency of the dynamics and upgrade the constructions to pathwise unique strong solutions of the ISDEs.
In Section 11, we establish uniqueness of the Dirichlet forms and prove the existence of the fully labeled diffusion process in $(\mathbb{R}^{d})^{\mathbb{N}}$.
In Section 12, we completes the program by proving convergence from finite-particle systems with reflecting boundary conditions to the infinite-particle ISDE dynamics.
Finally, the Appendix (Section 13) gathers auxiliary material on configuration spaces and Dirichlet forms, and recalls general ISDE theory from earlier works that is essential for our arguments.
Key difficulties and techniques. Two major obstacles underlie this program: the long-range and singular nature of Coulomb interactions. Previous methods relied on delicate hyperuniform estimates available only in special determinantal cases. Our approach replaces these with new analytic estimates for the logarithmic derivative, together with a quenched formulation that yields robust $L^{2}$ bounds. This combination provides a conceptually simple yet powerful tool, allowing us to establish strong well-posedness of Coulomb interacting Brownian motions in all dimensions $d\geq 2$ and for all $\beta>0$.
3. Concrete examples
We now present concrete examples of random point fields that satisfy assumptions ((A1))((A4)). Throughout, let $d\geq 2$ and $\beta>0$, and let $\Psi$ denote the $d$-dimensional Coulomb potential given in (1.1).
Example 3.1 ([1, 44]).
Let
[…]
where $c_{3.1}$ is a positive constant, and set
$\Phi^{N}(x)=N^{2/d}\,\Phi\!\left(\frac{x}{N^{1/d}}\right).$
Take $\Psi^{N}=\Psi$ and $O_{N}=\mathbb{R}^{d}$. Then ((A3)) is automatically satisfied.
In [44], a convergent subsequence was constructed, yielding the limiting Coulomb random point field $\mu$, which possesses bounded correlation functions of all orders together with other regularity properties. Moreover, the assumptions in [44] are weaker than those stated above; see [44, 1.25, Th. 6].
We now present two further specific examples of random point fields that satisfy assumptions ((A1))((A4)). These examples exhibit a form of “invariance” under certain transformations, which follows from the invariance of the corresponding finite-particle systems. As a consequence, the one-point correlation function is constant for each $N$-particle system.
Verifying that these examples satisfy ((A1))((A4)) requires a separate argument. The construction can also be extended to homogeneous spaces and, more generally, to compact Riemannian manifolds, including models beyond the periodic and spherical cases. These generalizations will be presented elsewhere.
Example 3.2 (Periodic approximation).
Let
$\mathbb{T}=(-\tfrac{1}{2},\tfrac{1}{2}]^{d}$
[…]
Set
$O_{N}=\Bigl{(}-\tfrac{N}{2},\,\tfrac{N}{2}\Bigr{)}^{d},\qquad\Psi^{N}({x},{y})=N^{2-d}\,\Psi_{\mathbb{T}}\!\left(\frac{{x}}{N},\,\frac{{y}}{N}\right).$
Note that, by construction, the measure $\mu$ is invariant under translations on $\mathbb{R}^{d}$.
@@ 3. Concrete examples
Let
$\mathbb{S}=\{\hat{x}\in\mathbb{R}^{d+1}:\lvert\hat{x}\rvert=1\}$
denote the unit sphere, and let $\mathbf{e}_{d+1}=(0,\dots,0,1)$ denote the north pole. Define the stereographic projection $\varpi:\mathbb{S}\setminus\{\mathbf{e}_{d+1}\}\to\mathbb{R}^{d}$ by
@@
$\varpi^{-1}(y)=\frac{(2y,\,|y|^{2}-1)}{1+|y|^{2}},\qquad y\in\mathbb{R}^{d},$
with Jacobian
$J(y)=\frac{2^{d}}{(1+|y|^{2})^{d}}.$
Set
$\Phi_{\mathbb{S}}({x})=1/\lvert J(\varpi^{-1}({x}))\rvert,\qquad\Phi^{N}({x})=N^{2-d}\,\Phi_{\mathbb{S}}({x}/N),$
and define
@@
where $\Delta_{\mathbb{S}}$ is the Laplace–Beltrami operator and $c_{3.3}$ is determined according to (2.15) and (2.16). Set
$\Psi^{N}({x},{y})=N^{2-d}\,\Psi_{\mathbb{S}}\!\left(\frac{{x}}{N},\,\frac{{y}}{N}\right),\qquad O_{N}=\mathbb{R}^{d}.$
Note that, by construction, the measure $\mu$ is invariant under the full Euclidean group $E(d)$, i.e., under all translations, rotations, and reflections.

### v4

@@ 2.8. Related works and positioning
We begin by recalling earlier developments on logarithmic derivatives, which constitute a foundational concept in the theory of infinite-particle dynamics.
Logarithmic derivatives. The concept of the logarithmic derivative for RPFs was introduced in [27].
In [27], it was proved that if $\nu$ admits a logarithmic derivative $\mathsf{d}^{\nu}$ and satisfies the quasi-Gibbs property, then (2.3) admits a weak solution, and the associated unlabeled dynamics $\mathsf{X}_{t}=\sum_{i}\delta_{X_{t}^{i}}$ is a $\nu$-reversible diffusion.
In [36], it was proved that, for general random point fields, the mere existence of a logarithmic derivative suffices to construct such dynamics.
However, the mere existence of a logarithmic derivative and the construction of a weak solution to the ISDE do not guarantee uniqueness of solutions or the existence of strong solutions. Moreover, they do not allow a detailed analysis of the qualitative behavior of the solutions. Addressing these issues is therefore essential.
Hyperuniformity. The logarithmic derivatives for the sineβ, Airyβ $(\beta=1,2,4)$, Bessel2, and Ginibre RPFs have been computed [27, 28, 35, 9]. The former three are one-dimensional systems with two-dimensional Coulomb interaction potentials, and hence are not Coulomb RPFs.
@@
These computations require strong hyperuniformity of particle numbers within $S_{R}$ for each $R\in\mathbb{N}$, a property that must be established separately for each RPF. For example, in the case of the Ginibre RPF, we use the following estimate [38, 32]:
@@
$\sup_{N\in\mathbb{N}}\mathrm{Var}^{\mu^{N}}\bigl[\mathsf{s}(S_{R})\bigr]=O(R)\quad\text{as }R\to\infty.$ (2.46)
@@
Such an estimate is called hyperuniform because, if $\mu^{N}$ is replaced by $\Lambda^{N}$, then the right-hand side becomes $O(R^{2})$ [23]. Here $\Lambda^{N}$ is the conditioned Poisson RPF defined by
$\Lambda^{N}=\Lambda\Bigl(\,\cdot\,\Big|\,\mathsf{s}(S_{R(N)})=N,\ \mathsf{s}(\mathbb{R}^{d}\backslash S_{R(N)})=0\Bigr).$
Here $\Lambda$ is the Poisson RPF on $\mathbb{R}^{d}$ with Lebesgue intensity $dx$, and $R(N)$ is the positive number such that the volume of $S_{R(N)}$ coincides with $N$.
While Leblé [23] established a hyperuniform estimate for general $\beta>0$, it is weaker than (2.46) and is insufficient for our purposes.
The proofs of the strong hyperuniform estimates for the above-mentioned examples rely on techniques specific to determinantal RPFs, and, except in the Ginibre case, do not extend to general Coulomb RPFs, which necessitates a new approach.
Logarithmic derivatives of Coulomb RPFs. Until now, the logarithmic derivatives of Coulomb RPFs have been computed only in the case of the two-dimensional Coulomb RPF at inverse temperature $\beta=2$, namely the Ginibre RPF [27].
In contrast, we present here a new, remarkably simple and transparent method that yields uniform estimates for the logarithmic derivative of $\mu^{N}$ (2.47) and for its conditional analogue (2.48), valid simultaneously for all spatial dimensions $d\geq 2$ and all inverse temperatures $\beta>0$.
The key insight is to exploit a certain harmonicity property of the interaction potential, which permits effective control over the infinite-dimensional interactions. Our approach is robust, conceptually clear, and provides a versatile tool for analyzing the stochastic dynamics of $d$-dimensional Coulomb RPFs, thereby overcoming the limitations of all previously known methods.
Annealed and quenched estimates for the logarithmic derivative. We introduce a robust method for estimating the logarithmic derivative of $\mu^{N}$ and refine it to conditional probabilities $\mu_{R,\mathsf{y}}^{N}$, exploiting a form of harmonicity of the interaction potential.
As a first step, we obtain the following uniform $L^{2}$-bound:
$\sup_{N\in\mathbb{N}}\int_{S_{Q}\times\mathsf{S}}\bigl|\mathsf{d}^{N}({x},\mathsf{s})\bigr|^{2}\,\mu^{N,[1]}(d{x}d\mathsf{s})<\infty.$ (2.47)
This bound provides the control needed to define logarithmic derivatives rigorously in the Coulomb setting (Lemma 4.3).
The proof of (2.47) is remarkably simple: it relies only on (2.9), (2.10), and the basic definition of the logarithmic derivative.
Moreover, the estimate (2.47) may be viewed as a form of hyperuniformity, since the quantity in (2.47) diverges when $\mu^{N,[1]}$ is replaced by the one-Campbell measure $\Lambda^{N,[1]}$ of the conditioned Poisson point process $\Lambda^{N}$, even after excluding a neighborhood of x{x}. This hyperuniformity is much weaker than that used in the previous works [16, 17, 18, 28, 27, 34, 35, 32], which motivates the development in this paper of a new framework for the analysis of stochastic dynamics associated with Coulomb random point fields.
We next establish the quenched analogue of (2.47):
$\sup_{R\geq Q+1}\sup_{N\in\mathbb{N}}\int_{S_{Q}\times\mathsf{S}}\bigl\lvert\mathsf{d}^{N}({x},\mathsf{s})\bigr\rvert^{2}\,\mu_{R,\mathsf{y}^{N}}^{N,[1]}(d{x}d\mathsf{s})<\infty\quad\text{for $\mu$-a.s.\,$\mathsf{y}$},$ (2.48)
where $\mathsf{y}^{N}$ is defined in (2.23) (Lemma 5.2).
Inequality (2.48) plays a decisive role in the proof of Theorem 2.1. Indeed, it fulfills the role previously played by strong hyperuniform estimates, but with a crucial difference: its proof is significantly simpler.
Earlier results relied on bounding averaged fluctuations of the exterior configuration, whereas here we directly control the external condition in a quenched form. This shift in perspective—replacing averaged estimates by a quenched analysis—is one of the key conceptual innovations of the present paper.
Interacting Brownian motions. Lang initiated the systematic study of interacting Brownian motions, which was subsequently developed by Fritz, Tanemura, and others in the case of finite-range interaction potentials [20, 21, 7, 42]. When the interaction becomes long-ranged, however, the problem of constructing solutions to ISDEs poses significant difficulties, even for potentials within Ruelle’s class—namely, those that are superstable and integrable at infinity. Later, the first author introduced the Dirichlet form approach, which made it possible to establish the existence of weak solutions to ISDEs for a broad family of interaction potentials [26, 27, 28, 29]. This framework encompasses, in particular, the Ginibre interacting Brownian motion. Nonetheless, the stronger questions of constructing pathwise unique strong solutions and establishing their uniqueness remained open.
In [46], Tsai succeeded in constructing pathwise unique strong solutions of ISDEs for the $\sin_{\beta}$ RPFs with all $\beta\geq 1$. These processes correspond to infinite-particle systems in one dimension with logarithmic interactions, that is, with the two-dimensional Coulomb potential. In this sense, the ISDE in [46] can be regarded as involving a Riesz potential. Tsai’s proof relies on a delicate coupling mechanism that exploits special properties of the logarithmic interaction in one dimension. Such an approach, however, has no counterpart in higher dimensions with genuine Coulomb potentials.
To apply this framework, one requires both an equilibrium state (RPF) and suitable regularity properties of the corresponding logarithmic derivative.
The construction of Coulomb RPFs was established in [1, 44]. By exploiting the quenched bound (2.48), we apply a Taylor expansion method adapted to the Coulomb interaction potential and prove the following explicit representation of the logarithmic derivative (Theorem 2.1):
$\mathsf{d}^{\mu}({x},\mathsf{s})=\beta\Big(-\nabla\Phi({x})+\sum_{s^{i}\in\overline{S}_{R}}\frac{{x}-s^{i}}{\lvert{x}-s^{i}\rvert^{d}}+\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{s}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{s}}^{\ell}({x})\Big).$
Using this representation, we further show that the associated ISDE satisfies the key assumption known as the IFC condition, which is required in the above theory (Theorem 2.3).
Further related works. Katori and Tanemura proposed an alternative approach to constructing stochastic dynamics with logarithmic interaction. Their method, based on space–time correlation functions, is restricted to the special case of inverse temperature $\beta=2$ in one-dimensional systems. For further details, see [12] and the references therein.
Moreover, [33] investigated in detail the relationship between the logarithmic derivative and the space–time correlation approach. It was rigorously established that the stochastic dynamics constructed by these two seemingly distinct methods are, in fact, identical. This coincidence highlights the fundamental role of the logarithmic derivative in unifying different perspectives on infinite-particle dynamics with logarithmic interactions.
In [2], Assiotis and Mirsajjadi constructed a non-equilibrium dynamics whose equilibrium distribution is the inverse Bessel RPF, and they identified the ISDE governing this dynamics. Their method is algebraic in nature, relying on intertwining theory and the use of derivatives of the so-called characteristic polynomial. This construction bears a striking resemblance to the framework of logarithmic derivatives developed in [27].
In [4], Bufetov and Kawamoto studied intertwining relations for Laguerre processes with parameter $\alpha$ across different particle numbers. They computed the logarithmic derivatives of the corresponding finite-particle systems and pointed out that the construction of the infinite-particle limit process remains an open problem, to be addressed in subsequent work.
In [14], Kawamoto constructed a unique strong solution to the ISDE associated with the generalized sine random point field at inverse temperature $\beta=2$, combining random matrix theory with the general ISDE theory [34].
In [6], Esaki and Tanemura constructed a unique strong solution to a jump-type ISDE associated with a jump-type infinite-particle system with long-range interaction, and constructed the interacting Cauchy processes.
In [40], Suzuki established the ergodicity of the unlabeled dynamics of the Ginibre interacting Brownian motion, thereby providing a concrete confirmation of the general ergodicity theory.
Key difficulties and techniques. A major obstacle underlying this program is the pronounced and extremely strong long-range interactions characteristic of Coulomb systems. Previous methods relied on delicate hyperuniform estimates, which are available only in the unique determinantal case $d=2$ and $\beta=2$.
Our approach replaces this reliance with new analytic estimates for the logarithmic derivative, together with a quenched formulation that yields robust $L^{2}$ bounds. This combination provides a conceptually simple yet powerful framework, which enables us to establish strong well-posedness of Coulomb interacting Brownian motions in all spatial dimensions $d\geq 2$ and for all inverse temperatures $\beta>0$.
Structure of the proofs. The proofs of the main theorems are distributed as follows: Theorem 2.1 is proved in Section 5, Theorem 2.2 in Section 9, Theorem 2.3 in Section 10, Theorem 2.4 in Section 11, and Theorem 2.5 in Section 12.
Sketch of the proof strategy. The main theorems are proved in Sections 412, and the overall strategy is as follows.
In Section 4, we establish the existence of the logarithmic derivative. A key ingredient is a robust uniform $L^{2}$-estimate for the logarithmic derivative, which exploits the harmonicity of the Coulomb potential.
Section 5 strengthens this result by proving a quenched estimate and deriving an explicit representation of the logarithmic derivative.
In Section 6, we establish non-collision properties of ISDE solutions by direct SDE techniques. These arguments do not rely on Dirichlet form theory; nevertheless, the resulting non-collision property is later used to prove uniqueness of Dirichlet forms.
Sections 79 form the technical core of the paper: we introduce Dirichlet form schemes for $m$-labeled dynamics and construct infinite-volume processes under both lower and upper Dirichlet forms.
In Section 10, using the explicit representation of the logarithmic derivatives, we verify the IFC condition ensuring consistency of the dynamics and upgrade the constructions to pathwise unique strong solutions of the ISDEs.
In Section 11, using the uniqueness of solutions of the ISDE, we establish uniqueness of extensions of the Dirichlet forms and prove the existence of the fully labeled diffusion process in $(\mathbb{R}^{d})^{\mathbb{N}}$.
In Section 12, we prove the convergence of finite-particle systems with reflecting boundary conditions using the convergence of the approximating lower Dirichlet forms. Finally, combining the uniqueness of extensions of Dirichlet forms with a sandwich argument, we prove that the $N$-particle dynamics converge to the infinite-particle ISDE dynamics.
3. Concrete examples
We present concrete examples of RPFs that satisfy assumptions ((A1))((A4)). Throughout this section, we assume $d\geq 2$ and $\beta>0$.
Example 3.1 ([44]).
Let $\Phi$ be such that $0\leq\Delta\Phi$ is bounded on $\mathbb{R}^{d}$, and set
$\Phi^{N}(x)=N^{2/d}\,\Phi\!\left(\frac{{x}+a_{N}}{N^{1/d}}\right).$
Here $a_{N}$ is a convergent sequence in $\mathbb{R}^{d}$ with limit $a$, and $a$ belongs to the droplet, that is, the support of the equilibrium measure associated with the potentials $\Phi$ and $\Psi$. Take $\Psi^{N}=\Psi$ and $O_{N}=\mathbb{R}^{d}$. Then ((A1)) ii and ((A3)) are automatically satisfied. Moreover, ((A1)) i is satisfied in a wide class of examples.
((A2)) i is obvious, and ii–iii follow from [44]. Indeed, in [44], a convergent subsequence was constructed, yielding the limiting Coulomb RPF $\mu$, which possesses bounded correlation functions of all orders, together with other regularity properties. Moreover, the assumptions in [44] are weaker than those stated above; see [44, 1.25, Th. 6].
If $d=2$ and $\Phi({x})=\frac{1}{2}\lvert x\rvert^{2}$, then the resulting ISDE is
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}-\frac{\beta}{2}\int_{0}^{t}X_{u}^{i}\,du-\frac{\beta}{2}at+\frac{\beta}{2}\int_{0}^{t}\lim_{R\to\infty}\sum_{\begin{subarray}{c}\lvert X_{u}^{j}\rvert\leq R\\ j\neq i\end{subarray}}\frac{X_{u}^{i}-X_{u}^{j}}{\lvert X_{u}^{i}-X_{u}^{j}\rvert^{2}}\,du.$
In this way, the term $\frac{\beta}{2}at$ appears, canceling the effect of the macroscopic position $a$ in the droplet.
We now present two further specific examples of RPFs that satisfy assumptions ((A1))((A4)). These examples exhibit a form of “invariance” under certain transformations, which follows from the invariance of the corresponding finite-particle systems. As a consequence, the one-point correlation function is constant for each $N$-particle system.
Example 3.2 (Periodic approximation).
@@ 3. Concrete examples
Let $\mathbb{S}=\{\hat{x}\in\mathbb{R}^{d+1}:\lvert\hat{x}\rvert=1\}$ denote the unit sphere, and let $\mathbf{e}_{d+1}=(0,\dots,0,1)$ denote the north pole. Define the stereographic projection $\varpi:\mathbb{S}\setminus\{\mathbf{e}_{d+1}\}\to\mathbb{R}^{d}$ by
@@
$\varpi^{-1}(x)=\frac{(2x_{1},\ldots,2x_{d},\,|x|^{2}-1)}{1+|x|^{2}},\qquad x\in\mathbb{R}^{d},$
@@
Define $O_{N}=\mathbb{R}^{d}$ and
$\Phi^{N}({x})=-N^{\frac{2-d}{d}}\,\log J(\varpi^{-1}({x}/N^{\frac{1}{d}})),$
$\Psi^{N}({x},{y})=N^{\frac{2-d}{d}}\,\Psi_{\mathbb{S}}\big(\varpi^{-1}({{x}}/{N^{\frac{1}{d}}}),\,\varpi^{-1}({{y}}/{N^{\frac{1}{d}}})\big).$
Note that, by construction, the measure $\mu$ is invariant under the full Euclidean group $E(d)$, i.e., under all translations, rotations, and reflections. The resulting diffusion $\mathbf{X}$ is invariant under the full Euclidean group $E(d)$.
Remark 3.1.
The construction of the Coulomb RPFs can be extended to homogeneous spaces and, more generally, to compact or weighted Riemannian manifolds, including models beyond the Gaussian, periodic, and spherical cases. These generalizations will be developed in future work.